1. Metallic silver can be prepared by the reaction of hydrogen gas on silver oxide.
  2. You react 16.7 g of Ag2O with 3.20 g H2?
  3. What is the limiting reagent?
  4. How many grams of Ag are produced?
  5. Arcylonitrile
1
Expert's answer
2022-01-25T10:08:02-0500

Ag2O + H2 = 2 Ag + H2 O .


Moles of Ag2O = 16.7 / 232 = 0.072 mol.

Moles of H2 = 3.2 / 2 = 1.6 mol.


Limiting reagent is Ag2O.

Moles of Ag Formed = ( 2 × 0.072 ) mol. = 0.144 mol.


Mass of Ag formed = 0.144 × 108 g

= 15.552 g Answer .


Acrylonitrile is a colorless, liquid, man-made chemical with a sharp, onion- or garlic-like odor. It can be dissolved in water and evaporates quickly. Acrylonitrile is used to make other chemicals such as plastics, synthetic rubber, and acrylic fibers.